A body thrown up vertically reaches a maximum height of 100m.Another body with double the mass thrown up with double the initial velocity reaches a maximum height h.what is the value of h?

Grade:Upto college level

2 Answers

Badiuddin askIITians.ismu Expert
14 years ago
Dear Alok 0 = U^2 - 2g* 100 U^2 = 2g*100 ... 1 for second mass 0 = (2U)^2 - 2g*h h = 4U^2 /2g from equation 1 h = 400 Please feel free to post as many doubts on our discussion forum as you can. If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly. We are all IITians and here to help you in your IIT JEE preparation. All the best. Regards, Askiitians Experts Badiuddin
Rithikaa
6 years ago
A body thrown vertically up with a velocity at a certain place reaches a maximum Height of 50 . Another body with double the mass thrown vertically up from the same place with half the initial velocity will reach a maximum height.